OpenStudy (anonymous):

will someone please teach me how to convert rectangular equations to polar and vice versa ... my test is 2mm and its already 2 AM ):

OpenStudy (dumbcow):

well there are 4 identities you will use: \[x^2 + y^2 = r^2\] \[x = r \cos \theta\] \[y =r \sin \theta\] \[\tan \theta = \frac{y}{x}\] the converting involves multiple substitutions

OpenStudy (dumbcow):

Here is example: \[y = 4x +3\] convert to polar \[r \sin \theta = 4r \cos \theta +3\] solve for r \[r = \frac{3}{\sin \theta - 4 \cos \theta}\]